The perimeter of a rectangle is 26 inches. The length of the rectangle is 1 more than three times the width. How many inches long are the width and the length of the rectangle?
Answer by TimothyLamb(4379)

---
2x + 2y = 26
x = 3y + 1
---
put the system of linear equations into standard form:
---
2x + 2y = 26
x - 3y = 1
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = length = 10 in
y = width = 3 in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php